OpenStudy (anonymous):

Last random question!! thanks to everyone that helped me 16/42,128

myininaya (myininaya):

?

OpenStudy (anonymous):

Division XD

OpenStudy (anonymous):

1/2633

myininaya (myininaya):

16/42=2(8)/[2(21)]=8/21

myininaya (myininaya):

so the comma is suppose to be division?

OpenStudy (anonymous):

like this? 

OpenStudy (anonymous):

oh no just 16 divided by 42,128

myininaya (myininaya):

\[\frac{8}{21} \div \frac{128}{1}=\frac{8}{21} \cdot \frac{1}{128}=\frac{8}{21} \cdot \frac{1}{8 \cdot 16}=\frac{1}{16 \cdot 21}\]

myininaya (myininaya):

what?

OpenStudy (anonymous):